Write the standard form of the equation of the circle with the given center (-2,-4) and radius of 6
Answer by Theo(13342)

standard form of equation of a circle is:
(x-h)^2 + (x-k)^2 = r^2
(h,k) is the center of the circle.
r is the radius.
in your equation, this becomes:
(x+2)^2 + (y+4)^2 = 36
to graph your equation, solve for y to get:
y = +/- sqrt(36 - (x+2)^2) - 4
that graph is shown below: